How do I solve this problem? Could you please show all work. We are working on interpolating and I need help, thanks in advance you guys are great. I was given the following table and problem.
Monthly life insurance premiums per 10,000 coverage for female non smokers
age / monthly life insurance premium per $10,000 coverage
20 / $2.00
30 / $3.16
40 / $4.32
50 / $5.48
60 / $6.64
If the premium is $6.90 for $20,000 coverage, find the womans age

Answer by stanbon(75887)

Every 10 years the premium goes up $1.16
$6.90 is 26 cents higher than the premium for a women
who is 60 years old.
Set up a proportion to see
26/116 = x/10
x=260/116=2.24 years
The woman is 2.24 or approximately 2 years and 3 month older that 60.
Cheers,
Stan H.